If I add four times my age four years from now to five times my age five years from now, I get ten times my current age. How old will I be five years from now?
Answer by mathitutor(25)

Assume my current age = x
4(x+4) + 5(x+5) = 10x
=> 4x + 16 + 5x + 25 = 10x
=> 9x + 41 = 10x
=> 10x-9x = 41
=> x = 41